OpenStudy (anonymous):

HELP :) Determine the x intercepts of the quadratic function f(x)= x^2+12x+37

OpenStudy (anonymous):

Looking at this equation, we can notice something very troublesome with trying to factor this guy. 37 is a prime number, only divisible by itself and 1. In order to solve this guy, you will need to use a technique called completing the square. After that, the function should be easier to factor, then you can determine the x intercepts.

OpenStudy (anonymous):

CAN YOU PLEASE HELP ME? :)

OpenStudy (anonymous):

complete the square i mean :) i suck at it lol

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrquad.htm ^this shows a good method for using the method I described. Because we want to find the x intercepts, we set y(x)=0 to get 0=x^2+12x+37 First take 37 and put it on the other side of the equation to get: -37=x^2+12x Then, divide by the coefficient in front of the x^2 term to get -37=x^2+12x -I know this is the same eq as above, it is just part of the process Now, take half of the coefficient in front of the x term and square it 12/2=6 6^2=36 (be careful not to forget the original sign of this term!!) Add the term from ^ to each side of the original equation to get -37+36= x^2+12x+36 Factor: -1=(x+6)^2 Normally, here we would take the square root and solve for x, however we cannot take the square root of a negative number. Thus, this function has no x intercepts. If you have access to a graphing calculator or software, try plugging in this function and graphing it. You will find that this graph does not intersect the x axis at any point

OpenStudy (anonymous):

This one is tricky because the question implies that there ARE x intercepts, however this cannot be true due to the computations above

OpenStudy (anonymous):

So there no intercepts? :)

OpenStudy (anonymous):

hey just use quadratic formula :)

OpenStudy (anonymous):

Yes, no intercepts. @some_someone, if you tried to do so with this equation you would find yourself trying to take the square root of a negative, just like we found ourselves doing in the above computations.

OpenStudy (anonymous):

exactly, you get complex/imaginary solutions. There is no actual real solutions :D

OpenStudy (anonymous):

lol, I don't think that was the focus of this exercise. However we basically learned HOW to complete the square and found something interesting about this function. What math are you taking @HopelessMathStudent ?

OpenStudy (anonymous):

btw, the complex solution would be (+/-)i-6=x, just in case you need to know the imaginary solution

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

You wouldn't know that if you didn't complete the square ;) lol

OpenStudy (anonymous):

haha i did the quadratic formula method :P

OpenStudy (anonymous):

Both methods are useful :D just depends on what you are comfortable with and what will lead to the correct(and sensible) solution

OpenStudy (anonymous):

agreed